OpenStudy (anonymous):

(3x+2)(2x^2-7x+3)=0... Solve the equation.

OpenStudy (anonymous):

(3x+2)(2x-1)(x-3)=0 x=-2/3, 1/2, 3

OpenStudy (anonymous):

oh sorry i put the wrong problem.

OpenStudy (anonymous):

\[(3x+4)(5x^2-14x+8)=0\]